Shift report 20000828 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Aug 28 2000 - 18:54:37 EDT

  • Next message: http account: "Shift report 20000829 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Aug 28 18:54:37 2000, Shift: 16:00-24:00
    Supervisor: fv, with: fv,kh,db
    Log book#5 pages 39 - 39
    
    Shift Summary:
    ------------------------------------------------------------
    AT the 4 o'clock meeting RHIC machine people decided to
    remove the IPM in 2 o'clock. They started access at 5 o'clock.
    The estimate is beam might be back around midnight, and with
    a reasonable vacuum back at the BRAHMS IR by tomorrow morning.
    This is potential good news.
    
    The brahms datataking is suspended until tomorrow morning.
    FR/DB will come in earlier than the regular day shift to 
    review the status and startup the runs.
    
    DB got a 'magnetic wand' from AGS. After having restored the
    control computer the D1,D2 and D5 field was turned to their nominal
    1/5 setting B. The response was identical in all 3 gaps.
    
    An updated shift list is available on the Web.
    
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Aug 28 2000 - 18:58:12 EDT